OpenStudy (anonymous):

In right triangle ABC, ∠A and ∠B are complementary angles and sin A = 89. What is cos B?

OpenStudy (welshfella):

there's something wrong with this question. the sine of an angle must be between -1 and 1 (89 is definitely not a sine!) Perhaps its A = 89 ? or sin A = 0.89?

OpenStudy (anonymous):

idk why it came out like that it a = 8/9

OpenStudy (welshfella):

oh ok 8/9 makes sense

OpenStudy (anonymous):

yeah idk why it came out as 89 when i pasted it

OpenStudy (welshfella):



OpenStudy (welshfella):

sin A = opposite side / hypotenuse = 8 /9 cos B = adjacent side / hypotenuse What is the adjacent side for angle B?

OpenStudy (welshfella):

note the hypotenuse is the longest (slanting) side

OpenStudy (welshfella):

complementary means <A and <B add up to 90 degrees that's why i could draw the right angled triangle